# Semantic-Aware Power Allocation for Generative Semantic Communications with Foundation Models

### v1

@@ Abstract
Recent advancements in diffusion models have made a significant breakthrough in generative modeling. The combination of the generative model and semantic communication (SemCom) enables high-fidelity semantic information exchange at ultra-low rates. A novel generative SemCom framework for image tasks is proposed, wherein pre-trained foundation models serve as semantic encoders and decoders for semantic feature extractions and image regenerations, respectively. The mathematical relationship between the transmission reliability and the perceptual quality of the regenerated image and the semantic values of semantic features are modeled, which are obtained by conducting numerical simulations on the Kodak dataset. We also investigate the semantic-aware power allocation problem, with the objective of minimizing the total power consumption while guaranteeing semantic performance. To solve this problem, two semantic-aware power allocation methods are proposed by constraint decoupling and bisection search, respectively. Numerical results show that the proposed semantic-aware methods demonstrate superior performance compared to the conventional one in terms of total power consumption.
@@ I Introduction
The end-to-end architecture is widely used in jointly training the neural network (NN) based semantic encoder and decoder, facilitating the formation and sharing of the knowledge base between them. The concept of deep joint source and channel coding (JSCC) was first proposed for image tasks by adopting the auto-encoder NN network [3], and numerous variants of deep JSCC were developed subsequently for various types of sources and channel models [4, 5]. These deep JSCC approaches have demonstrated superior performance over the conventional separated source and channel coding schemes in terms of distortion metrics such as mean square error (MSE), peak-signal-to-noise (PSNR), and multi-scale structural similarity (MS-SSIM). However, the distortion may no longer serve as the primary performance indicator for emerging applications with inference goals, where precisely conveying the semantic information becomes more important. To reserve the semantic, the authors in [6] proposed to integrate the generative adversarial network (GAN) into SemCom systems for signal regeneration. It was shown to significantly outperform the Deep JSCC technique in terms of both distortion and perceptual quality. Recent advancements in state-of-the-art diffusion models have marked a significant breakthrough in generative modeling, showing impressive results in regenerating images [7], audios [8], and videos [9]. The diffusion model has been adopted in SemCom systems for synthesizing semantic-consistent signals, utilizing a combination loss function of the MSE and Kullback-Leibler (KL) divergence [10]. This approach demonstrated high robustness to poor channel conditions and outperformed existing methods in generating high-quality images while preserving semantic information.
@@
Inspired by these, we propose the generative SemCom framework for image tasks by utilizing powerful pre-trained foundation models to extract semantic features and regenerate signals at the encoder and decoder, respectively. Within this framework, transmission reliability becomes the sole factor influencing the perceptual quality of the regenerated images, with their mathematical relationship modeled as a non-decreasing percpetion-error function. Semantic values of semantic data streams are defined to measure the semantic information accordingly. We investigate the semantic-aware resource allocation problem in the channel-uncoded case, aiming at minimizing the total power consumption while ensuring the perceptual quality of regenerated images. The rest of this paper is organized as follows. Section II introduces the proposed generated SemCom framework for image tasks and defines semantic values. Section III provides the semantic-aware power allocation problem formulation, and Section IV presents the proposed methods. Numerical results are given in Section V to demonstrate the performance of the proposed framework. Finally, Section V concludes this paper.
@@ II Generative SemCom Framework
The proposed generative SemCom framework for image task, as depicted in Fig. 1, consists of semantic encoder $\mathcal{F}_{en}$, transmission scheme $\mathcal{T}$, and semantic decoder $\mathcal{F}_{de}$. Before giving a detailed description of the generative SemCom framework, we introduce the semantic metric based on contrastive language-image pre-training (CLIP) similarity [12] to evaluate the perceptual quality of the regenerated image, which is written as
$P\triangleq\mathbb{E}\left[\mathrm{CLIP}\left(\mathbf{X},\hat{\mathbf{X}}\right)\right]=1-\mathbb{E}\left[\frac{F_{\mathrm{clip}}\left(\mathbf{X}\right)\cdot F_{\mathrm{clip}}\left(\hat{\mathbf{X}}\right)}{\left\|F_{\mathrm{clip}}\left(\mathbf{X}\right)\right\|\left\|F_{\mathrm{clip}}\left(\hat{\mathbf{X}}\right)\right\|}\right],$ (1)
where $P$ is within the range $[0,1]$. $\mathbf{X}$ and $\hat{\mathbf{X}}$ denote the source and the regenerated images, respectively. $F_{\mathrm{clip}}\left(\cdot\right)$ refers to a pre-trained model trained on a large text-image dataset [12].
@@ II-A Semantic Encoder
$\mathbf{S}_{i}=F_{en,i}\left(\mathbf{X}\mid\bm{\theta}_{i}^{*}\right),$ (2)
@@
To ensure compatibility with existing digital communication systems, the semantic feature $\mathbf{S}_{i}$ is converted into the bit sequence denoted as $\mathbf{K}_{i}$. We have $\mathbf{K}_{i}=\mathrm{\mathcal{B}}\left(\mathbf{S}_{i}\right)$, where $\mathrm{\mathcal{B}}\left(\cdot\right)$ is a binary mapping function such as ASCII, Unicode encoding and quantization. In SemCom systems, the semantic data streams contribute unequally to the perceptual quality of the regenerated image, which can be measured by the semantic metric closely related to the inference goal or task at the receiver. This is fundamentally different from conventional communication systems. Denote the semantic value of the $i$th semantic data stream as $L_{i}$ to quantify its semantic information in terms of a specific semantic metric. Generally, the semantic data stream with a larger $L_{i}$ has a greater impact on the perpetual quality of the regenerated signal, indicating its greater importance.
@@ II-C Semantic Decoder
In the semantic decoder, the pre-trained generative foundation model $F_{de}$, i.e., ControlNet [15] built upon the Stable Diffusion model [7] is employed to synthesize the received semantic data streams into an image $\hat{\mathbf{X}}$. In the channel-uncoded case, the received data streams regardless of transmission errors are processed by the generative foundation for signal synthesizing, as the transmission errors cannot be identified and corrected. The received semantic data streams $\hat{\mathbf{K}}_{i}$ are first reconverted into the semantic features $\hat{\mathbf{S}}_{i}=\mathcal{B}^{-1}\left(\hat{\mathbf{K}}_{i}\right)$ with $\mathrm{\mathcal{B}^{-1}}\left(\cdot\right)$ being the inverse operation of $\mathrm{\mathcal{B}}\left(\cdot\right)$. $\hat{\mathbf{S}}_{i}$ are forwarded to the generative foundation model $F_{de}$ for synthesizing $\hat{\mathbf{X}}$, which can be expressed as
@@
where $\bm{\omega}^{*}$ are the NN parameters of the ControlNet. Denote $\hat{L}_{i}$ as the semantic values of the $i$th received semantic data stream $\hat{\mathbf{K}}_{i}$. The semantic information is lossy due to the transmission errors, thus we have $\hat{L}_{i}\leq L_{i}$.
Given the semantic encoder and decoder, the transmission scheme and wireless channels remain to influence the perceptual quality of the regenerated image. As a consequence, the transmission reliability becomes the factor impacting the achieved perceptual quality. Denoting the bit error rate (BER) of the $j$th bit of $\hat{\mathbf{K}}_{i}$ as $\psi_{ij}$, the perception value $P$ defined in (1) becomes a function of $\psi_{ij}$.
@@
Assume that the perception value $P$ is non-decreasing with respective to (w.r.t.) the BER $\psi_{ij}$.
The semantic values of $i$th transmitted semantic data stream $\mathbf{K}_{i}$ is defined as
@@
III Problem Formulations of Semantic-aware Power Allocation
The transmission reliability significantly affects the perceptual quality of the regenerated images and the consumption of the resources. In contrast to conventional communications that treat the transmitted data streams equally, SemCom systems offer the opportunity to exploit the semantic importance to enhance resource efficiency. In this paper, we investigate the semantic-aware power allocation problem for generative SemCom systems at ultra-low rates. The objective is to minimize total power consumption while guaranteeing semantic performance.
@@ III Problem Formulations of Semantic-aware Power Allocation
where $h_{i}$ is channel assumed to be quasi-static and modelled as $h_{i}=\sqrt{h_{0}\left(\frac{d}{d_{0}}\right)^{-\alpha}}\tilde{h}_{i}$ where $h_{0}\left(\frac{d}{d_{0}}\right)^{-\alpha}$ is the path loss at distance $d$ with $h_{0}$ being the path loss at reference distance $d_{0}$. $\tilde{h}_{i}$ and $n_{i}$ are the Rayleigh fading channel with a covariance of $1$ and the Gaussian noise following the distributions of $n_{i}\sim\mathcal{CN}\left(0,\sigma_{i}^{2}\right)$. $q_{i}$ is the allocated power for each symbol of the $i$th semantic data stream.
@@
The optimal solutions $q_{i}^{*}$ to problem $\mathcal{P}1$ satisfies the equality of constraint (10b).
@@ IV-A Semantic-Aware Proportional Method
By assuming the independence of semantic data streams, the constraint (10b) can be decoupled into $I$ independent constraints, each corresponding to the semantic value constraint of an individual received data stream. Problem ($\mathcal{P}1$) is then relaxed into
@@
where $\bar{L}_{i}$ is the semantic value requirement of the $i$th received semantic data stream corresponding to $\bar{P}$. Based on Assumption 1, the semantic value of the received semantic data stream is non-increasing w.r.t. the BER $\psi_{i}$. Therefore, the optimal solutions to $\mathcal{P}2$ are obtained when the equalities of constraints (11b) hold. Denoting $\psi_{i}^{*}$ as the solution obtained by solving equation $\hat{L}_{i}\left(\psi_{i}\right)=\bar{L}_{i}$, the optimal solutions can be readily obtained by substituting $\psi_{i}^{*}$ back to (9), which is given by
@@ IV-B Semantic-Aware Bisection Method
Algorithm 1 Semantic-aware bisection method for two semantic extractors encoder
@@
$(\mathcal{P}3):\quad\quad\min_{\psi_{1},\psi_{2}}\quad$ $\sum_{i=1}^{2}\frac{K_{i}\sigma_{i}^{2}}{2|h_{i}|^{2}}\left(Q^{-1}\left(\psi_{i}\right)\right)^{2}$ (13a)
@@ V Numerical Results
To transmit the textual prompt and edge map data streams, the communication parameters are set as follows. The modulations of these two semantic data streams are the same. Both 8-QAM and 16-QAM modulation schemes are considered. The channel parameters are set to $d=100$ m, $d_{0}=1$ m, $h_{0}=-30$ dB and $\alpha=-3.4$. The noise power is set to $\sigma_{i}^{2}=-110$ dBm.
@@
Fig. 2 depicts two regenerated image examples using the proposed generative SemCom framework to demonstrate the achieved perceptual quality. As the increase of the BERs, the semantic performance in terms of the CLIP metric degrades. The compression rates achieved are $0.0278$ and $0.02597$ bits per pixel (BPP), indicating that ultra-low rates can be achieved within the proposed generative SemCom. It is difficult to explicitly obtain the mathematical relationship between the BERs and the perceptual quality of the regenerated image. Instead, we conduct numerical simulations on the Kodak dataset [17] to empirically derive this function. As shown in Fig. 3, the perception-error function is non-decreasing with BERs $\psi_{i}$, which is obtained by curve fitting using the numerical simulation points. Fig. 4 depicts the defined semantic values of both transmitted and received semantic data streams. The semantic values of textual prompt and edge map streams are $L_{1}=0.5887$ and $L_{2}=0.3596$, respectively. For the received semantic data streams, their semantic values, i.e., $\hat{L}_{1}$ and $\hat{L}_{2}$, are non-increasing with BERs $\psi_{i}$. In addition, the prompt feature has a greater impact on the CLIP performance compared to the edge map feature. However, the edge map feature exhibits greater vulnerability to the BER than the prompt feature due to its larger data length.
@@
The proposed semantic-aware proportional and bisection methods are compared with the conventional semantic-unaware one that treats the semantic data streams equally. For the semantic-unaware method, the SNRs for both semantic data streams are the same. For the semantic-proportional method, the allocated power is obtained based on (12), where $\frac{\hat{L}_{1}}{L_{1}}=\frac{\hat{L}_{2}}{L_{2}}$. The total power consumption comparison results are given in Fig. 5, showing that the total power consumption decreases as the increase of the performance requirement $\bar{P}$. Under stringent semantic performance requirements, the semantic-proportional method consumes lower power than the conventional approach. However, this performance advantage diminishes as $\bar{P}$ increases. The proposed semantic-aware bisection method consistently outperforms the semantic-aware proportional and the semantic-unaware methods. Moreover, it can be observed that higher modulation orders lead to increased power consumption due to lower transmission reliability. Notably, the performance advantage of the proposed semantic-aware methods over the semantic-unaware one becomes more evident as the increase of modulation order.

### v2

@@ Abstract
Recent advancements in diffusion models have led to a significant breakthrough in generative modeling. The combination of the generative model and semantic communication (SemCom) enables high-fidelity semantic information exchange at ultra-low rates. In this paper, a novel generative SemCom framework for image tasks is proposed, utilizing pre-trained foundation models as semantic encoders and decoders for semantic feature extraction and image regeneration, respectively. The mathematical relationship between transmission reliability and the perceptual quality of regenerated images is modeled and the semantic values of extracted features are defined accordingly. This relationship is derived through numerical simulations on the Kodak dataset. Furthermore, we investigate the semantic-aware power allocation problem, aiming to minimize total power consumption while guaranteeing semantic performance. To solve this problem, two semantic-aware power allocation methods are proposed by constraint decoupling and bisection search, respectively. Numerical results demonstrate that the proposed semantic-aware methods outperform conventional approach in terms of total power consumption.
@@ I Introduction
The end-to-end architecture is widely used to jointly train the neural network (NN)-based semantic encoder and decoder, facilitating the formation and sharing of the knowledge base between them. The concept of deep joint source and channel coding (JSCC) was first proposed for image tasks by adopting the auto-encoder NN network [3], and its numerous variants were developed subsequently to account for various types of sources and channel models [4, 5]. These deep JSCC approaches have demonstrated superior performance over the conventional separated source and channel coding schemes in terms of distortion metrics such as mean square error (MSE), peak-signal-to-noise (PSNR), and multi-scale structural similarity (MS-SSIM). However, the distortion may no longer serve as the primary performance indicator for emerging applications with inference goals, where precisely conveying the semantic information is sufficient and more important. To preserve semantic information, the authors in [6] proposed to integrate the generative adversarial network (GAN) into a SemCom systems for signal regeneration. This approach significantly outperformed the deep JSCC technique in terms of both distortion and perceptual quality. Recent advancements in state-of-the-art diffusion models have marked a significant breakthrough in generative modeling, demonstrating impressive results in regenerating images [7], audios [8], and videos [9]. The diffusion model has been adopted in SemCom systems for synthesizing semantic-consistent signals, utilizing a loss function combining both MSE and Kullback-Leibler (KL) divergence [10]. This approach demonstrated high robustness to poor channel conditions and outperformed existing methods in generating high-quality images while preserving semantic information.
@@
Inspired by these, we propose the generative SemCom framework for image tasks by utilizing powerful pre-trained foundation models to extract semantic features and regenerate signals at the encoder and decoder, respectively. Within this framework, transmission reliability becomes the sole factor influencing the perceptual quality of the regenerated images, with their mathematical relationship modeled as a non-decreasing percpetion-error function. The semantic values of semantic data streams are defined to measure the semantic information accordingly. We investigate the semantic-aware resource allocation problem in the channel-uncoded case, aiming to minimize the total power consumption while ensuring the perceptual quality of regenerated images. The rest of this paper is organized as follows. Section II introduces the proposed generated SemCom framework for image tasks and defines semantic values. Section III provides the semantic-aware power allocation problem formulation, and Section IV presents the proposed methods. Numerical results are given in Section V to demonstrate the performance of the proposed framework. Finally, Section V concludes this paper.
@@ II Generative SemCom Framework
The proposed generative SemCom framework for image tasks, as depicted in Fig. 1, consists of semantic encoder $\mathcal{F}_{en}$, transmission scheme $\mathcal{T}$, and semantic decoder $\mathcal{F}_{de}$. Before giving a detailed description of the generative SemCom framework, we introduce the semantic metric based on contrastive language-image pre-training (CLIP) similarity [12] to evaluate the perceptual quality of the regenerated image. The metric is written as
$P\triangleq\mathbb{E}\left[\mathrm{CLIP}\left(\mathbf{X},\hat{\mathbf{X}}\right)\right]=1-\mathbb{E}\left[\frac{F_{\mathrm{clip}}\left(\mathbf{X}\right)\cdot F_{\mathrm{clip}}\left(\hat{\mathbf{X}}\right)}{\big{\|}F_{\mathrm{clip}}\left(\mathbf{X}\right)\big{\|}\big{\|}F_{\mathrm{clip}}\left(\hat{\mathbf{X}}\right)\big{\|}}\right],$ (1)
where ${\mathbf{X}}$ and $\hat{\mathbf{X}}$ denote the source and the regenerated images, respectively. $F_{\mathrm{clip}}\left(\cdot\right)$ refers to a pre-trained model trained on a large text-image dataset [12].
@@ II-A Semantic Encoder
$\mathbf{S}_{i}=F_{en,i}\left(\mathbf{X}\mid\bm{\theta}_{i}^{*}\right),\forall i\in\{1,2\}$ (2)
@@
To ensure compatibility with existing digital communication systems, the semantic feature $\mathbf{S}_{i}$ is converted into the bit sequence denoted as $\mathbf{K}_{i}$. We have $\mathbf{K}_{i}=\mathrm{\mathcal{B}}\left(\mathbf{S}_{i}\right)$, where $\mathrm{\mathcal{B}}\left(\cdot\right)$ is a binary mapping function such as ASCII, Unicode encoding and quantization. In SemCom systems, the semantic data streams contribute unequally to the perceptual quality of the regenerated image. This varying contributions can be measured using a semantic metric closely related to the inference goal or task at the receiver. This is fundamentally different from conventional communication systems. Denote the semantic value of the $i$th semantic data stream as $L_{i}$ to quantify its semantic information in terms of a specific semantic metric. Generally, the semantic data stream with a larger $L_{i}$ has a greater impact on the perpetual quality of the regenerated signal, indicating its greater importance.
@@ II-C Semantic Decoder
In the semantic decoder, the pre-trained generative foundation model $F_{de}$, i.e., ControlNet [15] built upon the Stable Diffusion model [7], is employed to synthesize the received semantic data streams into an image $\hat{\mathbf{X}}$. In the channel-uncoded case, the received data streams regardless of transmission errors are processed by the generative foundation for signal synthesizing, as the transmission errors cannot be identified and corrected. The received semantic data streams $\hat{\mathbf{K}}_{i}$ are first reconverted into the semantic features $\hat{\mathbf{S}}_{i}=\mathcal{B}^{-1}\left(\hat{\mathbf{K}}_{i}\right)$ with $\mathrm{\mathcal{B}^{-1}}\left(\cdot\right)$ being the inverse operation of $\mathrm{\mathcal{B}}\left(\cdot\right)$. $\hat{\mathbf{S}}_{i}$ are forwarded to the generative foundation model $F_{de}$ to synthesize $\hat{\mathbf{X}}$, which can be expressed as
@@
where $\bm{\omega}^{*}$ are the NN parameters of the generative diffusion model. Denote $\hat{L}_{i}$ as the semantic values of the $i$th received semantic data stream $\hat{\mathbf{K}}_{i}$. The semantic information is lossy due to the transmission errors, thus we have $\hat{L}_{i}\leq L_{i}$.
Given the semantic encoder and decoder, the transmission scheme and wireless channels remain to influence the perceptual quality of the regenerated image. As a consequence, the transmission reliability becomes the sole factor impacting the achieved perceptual quality. Denoting the bit error rate (BER) of the $j$th bit of $\hat{\mathbf{K}}_{i}$ as $\psi_{ij}$, the perception value $P$ defined in (1) becomes a function of $\psi_{ij}$.
@@
Assume that the perception value $P$ is non-decreasing with respect to (w.r.t.) the BER $\psi_{ij}$.
The semantic value of $i$th transmitted semantic data stream $\mathbf{K}_{i}$ is defined as
@@
III Problem Formulation of Semantic-Aware Power Allocation
Transmission reliability significantly affects both the perceptual quality of the regenerated images and the resource consumption. Unlike conventional communication systems that treat all data streams equally, SemCom systems offer the opportunity to exploit the semantic importance to enhance resource efficiency. In this paper, we investigate the semantic-aware power allocation problem for the proposed generative SemCom framework. The objective is to minimize total power consumption while guaranteeing semantic performance.
@@ III Problem Formulation of Semantic-Aware Power Allocation
where $h_{i}$ is channel assumed to be quasi-static and modelled as $h_{i}=\sqrt{h_{0}\left(\frac{d}{d_{0}}\right)^{-\alpha}}\tilde{h}_{i}$ where $h_{0}\left(\frac{d}{d_{0}}\right)^{-\alpha}$ is the path loss at distance $d$ with $h_{0}$ being the path loss at reference distance $d_{0}$. $\tilde{h}_{i}$ is the Rayleigh fading channel with a covariance of $1$ and $n_{i}$ is the Gaussian noise following the distributions of $n_{i}\sim\mathcal{CN}\left(0,\sigma_{i}^{2}\right)$. $q_{i}$ is the allocated power for each symbol of the $i$th semantic data stream.
@@
The optimal solution $q_{i}^{*}$ to problem $\mathcal{P}1$ satisfies the equality of constraint (10b).
@@ IV-A Semantic-Aware Proportional Method
Assuming the independence of semantic data streams, the constraint (10b) can be decoupled into $I$ independent constraints, each corresponding to an individual semantic value constraint. Problem ($\mathcal{P}1$) is then relaxed into
@@
where $\bar{L}_{i}$ is the semantic value requirement of the $i$th received semantic data stream corresponding to the semantic performance requirement $\bar{P}$. Based on Assumption 1, the semantic value of the received semantic data stream is non-increasing w.r.t. the BER $\psi_{i}$. Therefore, the optimal solutions to $\mathcal{P}2$ are obtained when the equalities of constraints (11b) hold. Denoting $\psi_{i}^{*}$ as the solution obtained by solving equation $\hat{L}_{i}\left(\psi_{i}\right)=\bar{L}_{i}$, the optimal solutions can be readily obtained by substituting $\psi_{i}^{*}$ back to (9), which is given by
@@ IV-B Semantic-Aware Bisection Method
Algorithm 1 Semantic-aware bisection method
@@
$(\mathcal{P}3):\quad\quad\min_{\psi_{1},\psi_{2}}\quad$ $\sum_{i=1}^{2}\frac{K_{i}\sigma_{i}^{2}}{b_{i}|h_{i}|^{2}}\left(Q^{-1}\left(\frac{\log_{2}M_{i}}{a_{i}}\psi_{i}\right)\right)^{2}$ (13a)
@@ V Numerical Results
To transmit the textual prompt and edge map data streams, the communication parameters are set as follows. The modulations of these two semantic data streams are the same. Both 8-QAM and 16-QAM modulation schemes are considered. The channel parameters are set to $d=100$ m, $d_{0}=1$ m, $h_{0}=-30$ dB, and $\alpha=-3.4$. The noise power is set to $\sigma_{i}^{2}=-110$ dBm.
@@
Fig. 2 depicts two image examples using the proposed generative SemCom framework to demonstrate the achieved perceptual quality. As the increase of the BERs, the semantic performance in terms of the CLIP metric is degraded. The compression rates obtained are $0.0278$ and $0.02597$ bits per pixel (BPP), indicating that ultra-low rates can be achieved within the proposed generative SemCom framework. It is difficult to explicitly obtain the mathematical relationship between the BERs and the perceptual quality of the regenerated images. Instead, we conduct numerical simulations on the Kodak dataset [17] to empirically derive this function. As shown in Fig. 3, the perception-error function is non-decreasing with BERs $\psi_{i}$, which is obtained by curve fitting using the numerical simulation points. Fig. 4 depicts the defined semantic values of both transmitted and received semantic data streams. The semantic values of textual prompt and edge map streams are $L_{1}=0.5887$ and $L_{2}=0.3596$, respectively. For the received semantic data streams, their semantic values, i.e., $\hat{L}_{1}$ and $\hat{L}_{2}$, are non-increasing with BERs $\psi_{i}$. In addition, the prompt feature has a greater impact on the CLIP performance compared to the edge map feature. However, the edge map feature exhibits greater vulnerability to the BER than the prompt feature due to its larger data length.
@@
The proposed semantic-aware proportional and bisection methods are compared with the conventional semantic-unaware one that treats all data streams equally. For the semantic-unaware method, the SNRs for both semantic data streams are the same. For the semantic-proportional method, the allocated power is obtained based on (12), where $\frac{\hat{L}_{1}}{L_{1}}=\frac{\hat{L}_{2}}{L_{2}}$. The total power consumption comparison results are given in Fig. 5, showing that the total power consumption decreases as the increase of the performance requirement $\bar{P}$. Under stringent semantic performance requirements, the semantic-proportional method consumes lower power than the conventional approach. However, this performance advantage diminishes as $\bar{P}$ increases. The proposed semantic-aware bisection method consistently outperforms the semantic-aware proportional and the semantic-unaware methods. Moreover, it can be observed that higher modulation orders lead to increased power consumption due to lower transmission reliability. Notably, the performance advantage of the proposed semantic-aware methods over the semantic-unaware one becomes more evident as the increase of modulation order.